# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v2

@@ Abstract
The goal of alpha factor mining is to discover indicative signals of investment opportunities from the historical financial market data of assets, which can be used to predict asset returns and gain excess profits. Deep learning based alpha factor mining methods have shown to be powerful, which, however, lack of the interpretability, making them unacceptable in the risk-sensitive real markets. Alpha factors in formulaic forms are more interpretable and therefore favored by market participants, while the search space is complex and powerful explorative methods are urged. Recently, a promising framework is proposed for generating formulaic alpha factors using deep reinforcement learning, and quickly gained research focuses from both academia and industries. This paper first argues that the originally employed policy training method, i.e., Proximal Policy Optimization (PPO), faces several important issues in the context of alpha factors mining, making it ineffective to explore the search space of the formula. Herein, a novel reinforcement learning based on the well-known REINFORCE algorithm is proposed. Given that the underlying state transition function adheres to the Dirac distribution, the Markov Decision Process within this framework exhibit minimal environmental variability, making REINFORCE algorithm more appropriate than PPO. A new dedicated baseline is designed to theoretically reduce the commonly suffered high variance of REINFORCE. Moreover, the information ratio is introduced as a reward shaping mechanism to encourage the generation of steady alpha factors that can better adapt to changes in market volatility. Experimental evaluations on various real assets data show that the proposed algorithm can increase the correlation with asset returns by 3.83%, and a stronger ability to obtain excess returns compared to the latest alpha factors mining methods, which meets the theoretical results well.
@@ I Introduction
Constructing a superior portfolio involves distinguishing noise from the huge raw data in financial markets and discovering signals to balance risk and return, which is widely regarded as a signal processing problem[1, 2, 3]. In the realm of computational finance, it has become nearly a conventional practice to convert raw historical asset information into into alpha factor values[4], which serve as indicative signals of market trends for portfolio management[5]. The functions that generate these signals are known as alpha factors, which can be expressed in two distinct forms: the deep model and the formula[6, 7, 8]. Mining high-quality alpha factors has become a trendy topic among investors and researchers, due to their relationship with excess investment returns[9].
@@
Comparatively, alpha factors represented in formulaic forms enjoy much better interpretability and thus are favored by market participants. The most intuitive method for automatically mining formulaic factors is the tree-based models, which mutate expression trees to generate new alpha factors[11, 12, 13]. Another commonly used method is the genetic programming (GP), which generates the formulaic alpha factor expressions through iterative simulations of genetic variation and natural selection in biological evolution[14, 15, 16]. Unfortunately, both methods still face certain challenges. Tree-based models, while easy to understand and implement, may encounter performance bottlenecks when dealing with non-linear relationships and high-dimensional data[17]. Genetic programming, on the other hand, can deal with broader types of expressions by properly defining the search space, including complex nonlinear and non-parametric factor expressions, but it usually fails to explore the search space of large-scale expressions. Additionally, it is usually computationally expensive[18].
Recently, Yu et al. [19] proposes a promising framework named AlphaGen for mining formulaic alpha factors with Deep Reinforcement Learning (DRL), trying to bridge both deep learning and formulaic based methods. It employs Markov Decision Processes (MDPs)[20] to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using DRL. It essentially aims at finding formulaic alpha factors. With the help of DRL, it also overcomes the limitations of the explorative search ability suffered by traditional tree models and genetic programming[19].
@@
Based on the above discussions, this paper proposes a new DRL-based method for formulaic alpha factors mining, named QuantFactor REINFORCE (QFR). QFR abandons the critic network of AlphaGen and train the policy using the REINFORCE[25], a seminal Monte Carlo method for policy gradient estimation. In this regard, the biases of PPO can be vanished. The high variance of REINFORCE is mitigated in two aspects. For the variance caused by the environment, we observe that the underlying MDPs has deterministic transitions, which satisfies the Dirac distribution. That is, once a new action is selected, the new state is uniquely determined by the current sequence and the newly generated token within the RPN representation. This ensures that REINFORCE will not be degenerated by the environmental variance. For the variance caused by the policy gradient estimation, QFR employs a greedy policy to generate a novel baseline, which has been theoretically shown to bound the variance and leads to a lower variance than the original REINFORCE. Additionally, we introduce the Information Ratio (IR) for reward shaping to better balance returns and risks. As a result, QFR enjoys faster convergence speed to the optimal policy and produces steadier alpha factor signals across various market volatility.
To evaluate our proposed QFR algorithm, we conducted extensive experiments on multiple real-world asset datasets. Our experimental results show that the set of formulaic alpha factors generated by the QFR algorithm outperforms those generated by previous methods. It is also verified that theoretical results properly predict trends in the experimental results, even encountering distinct market volatility.
@@
  • We propose a more stable and efficient reinforcement learning algorithm for mining formulaic alpha factors. Unlike prior research that employs the actor-critic framework, we argue to discard the critic network due to the trajectory feedback and theoretically show that the high variance due to the absence of the critic network can be significantly mitigated.
  • Two new components are proposed for DRL-based formulaic alpha factors mining. A novel baseline that employs a greedy policy is proposed to reduce the variance of the policy gradient estimation. Additionally, IR is introduced as a reward shaping mechanism to balance returns and risks. We also provide a set of theoretical results, including an analysis of the training variance under state transition functions with various distributions, the derivation of the upper bound of the variance after introducing the baseline, and the proof that the variance decreases compared to REINFORCE.
  • Extensive experiments on multiple real-world asset datasets show that the proposed algorithm outperforms the reinforcement learning algorithm used in work [19], as well as various tree-based and heuristic algorithms. Compared to the previous state-of-the-art algorithm, it improves the correlation with asset returns by 3.83%, while also demonstrating a stronger ability to generate excess profits. Additionally, the experimental results closely align with the theoretical results.
The rest of this paper is organized as follows. Related work is briefly reviewed in Section II. The formulaic alpha factors for predicting asset prices are introduced and the corresponding MDPs is formulated in Section III. Section IV details the proposed QFR algorithm and provides thorough theoretical analyses. Numerical results are presented in Section V to show the performance of the QFR. Finally, Section VI concludes this paper.
@@ II-A Automatic Mining for Alpha Factors
Alpha factors are generally represented in the form of deep models or formulas. Alpha factors using end-to-end deep models are more complex and usually trained with supervised learning[26, 27, 4], utilizing MLP[28] or sequential models like LSTM[29] and Transformer[30] to extract features embedded in the historical data of assets. Recently, reinforcement learning has attracted much attention in the context of computational finance and fintech. It becomes a key technology in alpha factor mining[19], investment portfolio optimization[31], and risk management design[32], thanks to its advantages in handling non-linearity, high-dimensional data, and dynamic environments[33]. By modeling market characteristics as states, maker or taker orders as actions, and profit and loss as rewards, reinforcement learning can also be used to train deep policy models that represents alpha factors[34, 35, 36].
@@
On the other hand, alpha factors represented in formulaic forms have much better interpretability and thus are favored by market participants. In the past, these formulaic alpha factors were constructed by human experts using their domain knowledge and experience, often embodying clear economic principles. For example, Kakushadze [37] presents 101 formulaic alpha factors tested in the US asset market. However, the alpha factor mining process relying on human experts suffers from multiple drawbacks such as strong subjectivity[38], time-consuming[38], insufficient risk control[39], and high costs[40]. To address these issues, algorithms for automatically mining formulaic alpha factors have been proposed [19], such as tree models represented by GBDT[11], XGBoost[12], and LightGBM[13], as well as heuristic algorithms represented by GP[14, 15, 16]. These algorithms can quickly discover numerous new formulaic alpha factors without requiring the domain knowledge or experience of human experts. They offer performance comparable to more complex deep learning-based alpha factors while maintaining relatively high interpretability.
@@
The work in [19] is the first to employ DRL for formulaic alpha factors mining. It employs MDPs to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using reinforcement learning. Compared to works [34, 35, 36] that directly use reinforcement learning to build trading agents, this framework uses historical quantitative and price data of assets as input to find a set of formulaic alpha factors with strong interpretability, avoiding the black-box problem and overcoming the limitations of [11, 12, 13, 14, 15] in independently mining individual formulaic factors, such as homogenization among factors, lack of synergy, and difficulty in adapting to dynamic market changes. The characteristics of reinforcement learning, supervised learning, tree models, and heuristic algorithms when applied to mining alpha factors are shown in Table I.
@@ II-B The REINFORCE algorithms
Williams is the first to intodcuce the REINFORCE algorithm in his works[25]. The algorithm is straightforward and versatile, suitable for a wide range of tasks that can be modeled as MDPs[42]. However, in MDPs with stochastic state transitions and immediate rewards, it often underperforms compared to actor-critic methods[43]. Actor-critic methods, which utilizes value-based techniques to decrease variance and integrate temporal-difference learning for dynamic programming principles, are generally favored. Consequently, the REINFORCE algorithm has not gained widespread favor in the reinforcement learning community. However, our work demonstrates that the REINFORCE algorithm can be suitable to MDPs for mining formulaic factors as long as a proper baseline is used to reduce the variance.
A series of studies have explored the incorporation of a baseline value in the REINFORCE algorithm. To our best knowledge, [44] is the pioneer in demonstrating that employing the expected reward as a baseline does not universally decrease variance, and he introduced an optimal baseline for a basic 2-armed bandit scenario to address the issue. Subsequently, formal examinations of baseline usage were systematically conducted within the online learning context[45, 46]. Additionally, the regret of REINFORCE algorithm was studied[47]. Recently, The work in [48] delved into the intricacies of multi-armed bandit optimization, revealing that the expected, as opposed to the stochastic, gradient ascent in REINFORCE can lead to a globally optimal solution. Building on this, the work was extended to the function of the baseline value in the natural policy gradient[49], concluding that variance reduction is not critical for natural policy gradient [50]. More recently, the work in [24] find that the REINFORCE is suitable for reinforcement learning from human feedback in large language models, and proposed a baseline value serves as a kind of normalization by comparing the rewards of a random response with those of the greedy response.
@@ III Problem Formulation and Preliminaries
This section first introduces the definition of formulaic alpha factors and their sequences with RPN. Next, the MDPs for mining the formulaic alpha factors modeled by Yu et al. [19] are detailed. Lastly, the two seminal methods for solving the modeled MDPs, i.e., PPO and REINFORCE, are compared to motivate the proposed QFR algorithm.
@@ III-A Alpha Factors for Predicting Asset Prices
After the policy model outputs a new formulaic alpha factor, this alpha will be added to an alpha factors pool $\mathcal{F}$, say =$\mathcal{F}=\{f_{1},f_{2},...,f_{K}\}$ with $K$ factors. The linear combination model [51] is adopted to compute the asset price predicted by the alpha factor pool[19]. Specifically, suppose each $k$-th alpha factor ($1\leq k\leq K$) is associated with a weight $w_{k}$, and the asset price is calculated as $\mathbf{z}_{l}^{\prime}=\sum^{K}_{k=1}w_{k}f_{k}(\mathbf{X}_{l})$. In other words, the mining of alpha factors is to search a set of alpha factors and use them in combination. The weight of each factor indicates its exposure on the assets and is optimized using gradient descent. The loss function of learning the weights vector $\mathbf{\omega}\in\mathbb{R}^{K\times 1}$ is defined as the mean squared error (MSE) between the model output and the ground-truth asset prices $\mathcal{Y}=\{\mathbf{y}_{l}\}$ with $l\in\{1,2,\cdots,L\}$ and $\mathbf{y}_{l}\in\mathbb{R}^{n\times 1}$:
$L(\mathbf{\omega})=\frac{1}{L}\sum_{l=1}^{L}\left\|\mathbf{z}_{l}^{\prime}-\mathbf{y}_{l}\right\|^{2},$ (1)
Due to significant differences in the scales of different alpha factor values, they are first normalized to have a mean of 00 and a maximum value of $1$. If the number of factors in the factor pool exceeds a certain threshold, the alpha factor with the smallest weight and its corresponding weight will be discarded.
@@
TABLE II: Some Alpha Factor Examples from Alpha 101
Alpha101 Index Formulaic Expression RPN Representation
Alpha#6 Mul(-1, Corr(open, volume, 10d)) BEG -1 open volume 10d Corr Mul SEP
Alpha#12 Mul(Sign(Delta(volume, 1d)), Mul(-1, Delta(close, 1d))) BEG volume 1d Delta Sign -1 close 1d Delta Mul Mul SEP
Alpha#41 Div(Pow(Mul(high, low), 0.5), vwap) BEG high low Mul 0.5 Pow vwap Div SEP
Alpha#101 Div(Sub(close, open), Add(Sub(high, low), 0.001)) BEG close open Sub high low Sub 0.001 Add Div SEP
III-B Formulaic Alpha Factors
Formulaic alpha factors are mathematical expressions that can be represented using RPN, which is a sequence of tokens. Tokens include various operators, the original volume-price features, fundamental features, time deltas, constants, and sequence indicators. The operators include elementary functions that operate on single-day data, known as cross-sectional operators (e.g., Abs($x$) for the absolute value $|x|$, Log($x$) for the natural logarithm log($x$)), as well as functions that operate on a series of daily data, known as time-series operators (e.g., Ref($x,l$) for the expression $x$ evaluated at $l$ days before the current day, where $l$ denotes a time token, such as 10d (10 days)). The Begin (BEG) token and Separator (SEP) token of the RPN representation are used to mark the beginning and end of the sequence. Table III illustrates a selection of these tokens as examples. Such formulas can naturally be represented by an expression tree, with each non-leaf node representing an operator, and the children of a node representing the original volume-price features, fundamental features, time deltas, and constants being operated on. Each such expression has a unique post-order traversal, using RPN. An example of a formulaic alpha expression, together with its corresponding tree and RPN sequence, is shown in Fig. 1. More examples of formulaic alpha expressions derived from the well-known Alpha101 [37] and their corresponding RPN sequences are shown in Table II.
TABLE III: Examples of Formulaic Tokens
@@ III-B Formulaic Alpha Factors
where $\mathrm{Cov}(\cdot,\cdot)$ indicates the covariance matrix between two vectors, and $\sigma_{\cdot}$ means the standard deviation of a vector. The averaged IC values over all $L$ trading days are denoted as $\overline{IC}=\mathbb{E}_{l}\left[IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)\right]=\frac{1}{L}\sum^{L}_{l=1}IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)$.

### v3

@@ Abstract
Alpha factor mining aims to discover investment signals from the historical financial market data, which can be used to predict asset returns and gain excess profits. Powerful deep learning methods for alpha factor mining lack interpretability, making them unacceptable in the risk-sensitive real markets. Formulaic alpha factors are preferred for their interpretability, while the search space is complex and powerful explorative methods are urged. Recently, a promising framework is proposed for generating formulaic alpha factors using deep reinforcement learning, and quickly gained research focuses from both academia and industries. This paper first argues that the originally employed policy training method, i.e., Proximal Policy Optimization (PPO), faces several important issues in the context of alpha factors mining. Herein, a novel reinforcement learning algorithm based on the well-known REINFORCE algorithm is proposed. REINFORCE employs Monte Carlo sampling to estimate the policy gradient—yielding unbiased but high variance estimates. The minimal environmental variability inherent in the underlying state transition function, which adheres to the Dirac distribution, can help alleviate this high variance issue, making REINFORCE algorithm more appropriate than PPO. A new dedicated baseline is designed to theoretically reduce the commonly suffered high variance of REINFORCE. Moreover, the information ratio is introduced as a reward shaping mechanism to encourage the generation of steady alpha factors that can better adapt to changes in market volatility. Evaluations on real assets data indicate the proposed algorithm boosts correlation with returns by 3.83%, and a stronger ability to obtain excess returns compared to the latest alpha factors mining methods, which meets the theoretical results well.
@@ I Introduction
Constructing a superior portfolio involves distinguishing noise from the huge raw data in financial markets and discovering signals to balance risk and return, which is widely regarded as a signal processing problem[1, 2, 3]. In the realm of computational finance, it has become nearly a conventional practice to convert raw historical asset information into alpha factor values[4], which serve as indicative signals of market trends for portfolio management[5]. The functions that generate these signals are known as alpha factors, which can be expressed in two distinct forms: the deep model and the formula[5]. Mining high-quality alpha factors has become a trendy topic among investors and researchers, due to their relationship with excess investment returns[6].
@@
Comparatively, alpha factors represented in formulaic forms enjoy much better interpretability and thus are favored by market participants. The most intuitive method for automatically mining formulaic factors is the tree-based models, which mutate expression trees to generate new alpha factors[8, 9, 10]. Another commonly used method is the genetic programming (GP), which generates the formulaic alpha factor expressions through iterative simulations of genetic variation and natural selection in biological evolution[11]. Unfortunately, both methods still face certain challenges. Tree-based models, while easy to understand and implement, may encounter performance bottlenecks when dealing with non-linear relationships and high-dimensional data[12]. Genetic programming, on the other hand, can deal with broader types of expressions by properly defining the search space, including complex nonlinear and non-parametric factor expressions, but it usually fails to explore the search space of large-scale expressions. Additionally, it is usually computationally expensive[13].
Recently, Yu et al. [14] proposes a promising framework named AlphaGen for mining formulaic alpha factors with Reinforcement Learning (RL), trying to bridge both deep learning and formulaic based methods. It employs Markov Decision Processes (MDPs)[15] to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using RL. It essentially aims at finding formulaic alpha factors. With the help of RL, it also overcomes the limitations of the explorative search ability suffered by traditional tree models and genetic programming[14].
@@
Based on the above discussions, this paper proposes a new RL-based method for formulaic alpha factors mining, named QuantFactor REINFORCE (QFR). QFR abandons the critic network of AlphaGen and train the policy using the REINFORCE[20], a seminal Monte Carlo method for policy gradient estimation. In this regard, the biases of PPO can be vanished. The high variance of REINFORCE is mitigated in two aspects. For the variance caused by the environment, we observe that the underlying MDPs has deterministic transitions, which satisfies the Dirac distribution. That is, once a new action is selected, the new state is uniquely determined by the current sequence and the newly generated token within the RPN representation. This ensures that REINFORCE will not be degenerated by the environmental variance. For the variance caused by the policy gradient estimation, QFR employs a greedy policy to generate a novel baseline, which has been theoretically shown to bound the variance and leads to a lower variance than the original REINFORCE. Additionally, we introduce the Information Ratio (IR) for reward shaping to better balance returns and risks. As a result, QFR enjoys faster convergence speed to the optimal policy and produces steadier alpha factor signals across various market volatility.
To evaluate our proposed QFR algorithm, we conducted extensive experiments on multiple real-world asset datasets. Our experimental results show that the set of formulaic alpha factors generated by QFR outperforms those generated by previous methods. It is also verified that theoretical results properly predict trends in the experimental results, even encountering distinct market volatility.
@@
  • We propose a more stable and efficient RL algorithm for mining formulaic alpha factors. Unlike prior research that employs the actor-critic framework, we argue to discard the critic network due to the trajectory feedback and theoretically show that the high variance due to the absence of the critic network can be significantly mitigated.
  • Two new components are proposed for RL-based formulaic alpha factors mining. A novel baseline that employs a greedy policy is proposed to reduce the variance of the policy gradient estimation. Additionally, IR is introduced as a reward shaping mechanism to balance returns and risks. We also provide a set of theoretical results, including an analysis of the training variance under state transition functions with various distributions, the derivation of the upper bound of the variance after introducing the baseline, and the proof that the variance decreases compared to REINFORCE.
  • Extensive experiments on multiple real-world asset datasets show that the proposed algorithm outperforms the RL algorithm used in work [14], as well as various tree-based and heuristic algorithms. Compared to the previous state-of-the-art algorithm, it improves the correlation with asset returns by 3.83%, while also demonstrating a stronger ability to generate excess profits. Additionally, the experimental results closely align with the theoretical results.
The rest of this paper is organized as follows. The formulaic alpha factors for predicting asset prices are introduced and the corresponding MDPs is formulated in Section II. Section III details the proposed QFR algorithm and provides thorough theoretical analyses. Numerical results are presented in Section IV to show the performance of the QFR. Finally, Section V concludes this paper.
@@ II-A Automatic Mining for Alpha Factors
Alpha factors are generally represented in the form of deep models or formulas. Alpha factors using end-to-end deep models are more complex and usually trained with supervised learning[4], utilizing Multilayer Perceptron (MLP)[21, 22, 23, 24, 25] or sequential models like Long Short Term Memory (LSTM)[26] and Transformer[27] to extract features embedded in the historical data of assets. Recently, reinforcement learning has attracted much attention in the context of computational finance and fintech. It becomes a key technology in alpha factor mining[14], investment portfolio optimization[28], and risk management design[29], thanks to its advantages in handling non-linearity, high-dimensional data, and dynamic environments[30]. By modeling market characteristics as states, maker or taker orders as actions, and profit and loss as rewards, RL can also be used to train deep policy models that represents alpha factors[31, 32, 33].
@@
On the other hand, alpha factors represented in formulaic forms have much better interpretability and thus are favored by market participants. In the past, these formulaic alpha factors were constructed by human experts using their domain knowledge and experience, often embodying clear economic principles. For example, Kakushadze [34] presents 101 formulaic alpha factors tested in the US asset market. However, the alpha factor mining process relying on human experts suffers from multiple drawbacks such as strong subjectivity[35], time-consuming[35], insufficient risk control[36], and high costs[37]. To address these issues, algorithms for automatically mining formulaic alpha factors have been proposed [14], such as tree models represented by Gradient Boosting Decision Tree (GBDT)[8], eXtreme Gradient Boosting (XGBoost)[9], and Light Gradient Boosting Machine (LightGBM)[10], as well as heuristic algorithms represented by GP[38]. These algorithms can quickly discover numerous new formulaic alpha factors without requiring the domain knowledge or experience of human experts. They offer performance comparable to more complex deep learning-based alpha factors while maintaining relatively high interpretability.
@@
The work in [14] is the first to employ RL for formulaic alpha factors mining. It employs MDPs to simulate the generation process of formulaic alpha factors and trains policies to directly generate a set of collaborative formulaic alpha factors using RL. Compared to works [31, 32, 33] that directly use RL to build trading agents, this framework uses historical quantitative and price data of assets as input to find a set of formulaic alpha factors with strong interpretability, avoiding the black-box problem and overcoming the limitations of [8, 9, 10, 38] in independently mining individual formulaic factors, such as homogenization among factors, lack of synergy, and difficulty in adapting to dynamic market changes. The characteristics of RL, supervised learning, tree models, and heuristic algorithms when applied to mining alpha factors are shown in Table I.
@@ II-B The REINFORCE algorithms
Williams is the first to introduce the REINFORCE algorithm in his works[20]. The algorithm is straightforward and versatile, suitable for a wide range of tasks that can be modeled as MDPs[40]. However, in MDPs with stochastic state transitions and immediate rewards, it often underperforms compared to actor-critic methods[41]. Actor-critic methods, which utilizes value-based techniques to decrease variance and integrate temporal-difference learning for dynamic programming principles, are generally favored. Consequently, the REINFORCE algorithm has not gained widespread favor in the RL community. However, our work demonstrates that the REINFORCE algorithm can be suitable to MDPs for mining formulaic factors as long as a proper baseline is used to reduce the variance.
A series of studies have explored the incorporation of a baseline value in the REINFORCE algorithm. To our best knowledge, [42] is the pioneer in demonstrating that employing the expected reward as a baseline does not universally decrease variance, and he introduced an optimal baseline for a basic 2-armed bandit scenario to address the issue. Additionally, the regret of REINFORCE algorithm was studied[43]. Recently, The work in [44] delved into the intricacies of multi-armed bandit optimization, revealing that the expected, as opposed to the stochastic, gradient ascent in REINFORCE can lead to a globally optimal solution. Building on this, the work was extended to the function of the baseline value in the natural policy gradient[45], concluding that variance reduction is not critical for natural policy gradient [46]. More recently, the work in [19] find that the REINFORCE is suitable for reinforcement learning from human feedback in large language models, and proposed a baseline value serves as a kind of normalization by comparing the rewards of a random response with those of the greedy response.
@@ III Problem Formulation and Preliminaries
This section first introduces the definition of formulaic alpha factors and their RPN sequences. Next, the MDPs for mining formulaic alpha factors modeled by Yu et al. [14] are detailed. Lastly, the two seminal methods for solving the modeled MDPs, i.e., PPO and REINFORCE, are compared to motivate the proposed QFR algorithm.
@@ III-A Alpha Factors for Predicting Asset Prices
After the policy model outputs a new formulaic alpha factor, this alpha will be added to an alpha factors pool $\mathcal{F}$, say =$\mathcal{F}=\{f_{1},f_{2},...,f_{K}\}$ with $K$ factors. The linear combination model is adopted to compute the asset price predicted by the alpha factor pool[14]. Specifically, suppose each $k$-th alpha factor ($1\leq k\leq K$) is associated with a weight $w_{k}$, and the asset price is calculated as $\mathbf{z}_{l}^{\prime}=\sum^{K}_{k=1}w_{k}f_{k}(\mathbf{X}_{l})$. In other words, the mining of alpha factors is to search a set of alpha factors and use them in combination. The weight of each factor indicates its exposure on the assets and is optimized using gradient descent. The loss function of learning the weights vector $\mathbf{\omega}\in\mathbb{R}^{K\times 1}$ is defined as the mean squared error (MSE) between the model output and the ground-truth asset prices $\mathcal{Y}=\{\mathbf{y}_{l}\}$ with $l\in\{1,2,\cdots,L\}$ and $\mathbf{y}_{l}\in\mathbb{R}^{n\times 1}$:
$L(\mathbf{\omega})=\frac{1}{L}\sum_{l=1}^{L}\left\|\mathbf{z}_{l}^{\prime}-\mathbf{y}_{l}\right\|^{2}.$ (1)
Due to significant differences in the scales of different alpha factor values, they are first normalized to have a mean of 00 and a maximum value of $1$. If the number of factors in the pool exceeds a certain threshold, the alpha factor with the smallest weight and its corresponding weight will be discarded.
III-B Formulaic Alpha Factors
Formulaic alpha factors are mathematical expressions that can be represented using RPN, which is a sequence of tokens. Tokens include various operators, the original volume-price features, fundamental features, time deltas, constants, and sequence indicators. The operators include elementary functions that operate on single-day data, known as cross-sectional operators (e.g., Abs($x$) for the absolute value $|x|$, Log($x$) for the natural logarithm log($x$)), as well as functions that operate on a series of daily data, known as time-series operators (e.g., Ref($x,l$) for the expression $x$ evaluated at $l$ days before the current day, where $l$ denotes a time token, such as 10d (10 days)). The Begin (BEG) token and Separator (SEP) token of the RPN representation are used to mark the beginning and end of the sequence. Table II illustrates a selection of these tokens as examples. Such formulas can naturally be represented by an expression tree, with each non-leaf node representing an operator, and the children of a node representing the original volume-price features, fundamental features, time deltas, and constants being operated on. Each such expression has a unique post-order traversal, using RPN. An example of a formulaic alpha expression, together with its corresponding tree and RPN sequence, is shown in Fig. 1.
@@ III-B Formulaic Alpha Factors
TABLE II: Examples of Formulaic Tokens
@@
where $\mathrm{Cov}(\cdot,\cdot)$ indicates the covariance matrix between two vectors, and $\sigma_{\cdot}$ means the standard deviation of a vector. The IC serves as a critical metric in quantitative finance, where higher values indicate stronger predictive power. This directly translates to portfolio enhancement through more informed capital allocation: factors with elevated IC enable investors to overweight assets with higher expected returns while underweighting underperformers. The averaged IC values over all $L$ trading days are denoted as $\overline{IC}=\mathbb{E}_{l}\left[IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)\right]=\frac{1}{L}\sum^{L}_{l=1}IC\left(\mathbf{z}^{\prime}_{l},\mathbf{y}_{l}\right)$.